OpenStudy (gabyramirezzz):

Sales tax word problem? Algebra 2

OpenStudy (gabyramirezzz):

In the state of Connecticut there is a state sales tax of 6.35% on all goods purchased. If jan buys a TV for $599, what will be the final bill, including state sales tax?

Awolflover1 (awolflover1):

hmm....what do you think? \[599+599\times0.0635 ?? \]

OpenStudy (gabyramirezzz):

why would we add 599 twice?

Awolflover1 (awolflover1):

\[$599 + 6.35%=605.35 \]

Awolflover1 (awolflover1):

=605.35 right?

OpenStudy (gabyramirezzz):

right

Awolflover1 (awolflover1):

im little confused it does say '' including sales tax.''

OpenStudy (gabyramirezzz):

yes

Awolflover1 (awolflover1):

so 599 on both sides ? @agent0smith

OpenStudy (gabyramirezzz):

wait i think i just figured it out

OpenStudy (gabyramirezzz):

all you have to do is multiply 599 by 0.0635 which =38.04

OpenStudy (gabyramirezzz):

wait nevermind

Awolflover1 (awolflover1):

hmm @DanJS :\

OpenStudy (danjs):

yeah, Total cost = Item Cost + Sales Tax Total Cost = 599 + 599*0.0634 = $637.04 you calculated wrong to start

OpenStudy (gabyramirezzz):

oh my gosh that's what she was saying from the start!!

Awolflover1 (awolflover1):

dan saves the day :')

OpenStudy (gabyramirezzz):

THANK YOU!!

OpenStudy (danjs):

welcome